# Towards Verifiable Text Generation with Evolving Memory and Self-Reflection

### v2

@@ Abstract
Towards Verifiable Text Generation with Evolving Memory and Self-Reflection
Hao Sun1, Hengyi Cai2, Bo Wang4, Yingyan Hou2 Xiaochi Wei3, Shuaiqiang Wang3, Yan Zhang1, Dawei Yin3 1Peking University,2Chinese Academy of Sciences 3Baidu Inc, 4Beijing Institute of Technology sunhao@stu.pku.edu.cn
@@ 1 Introduction
However, verifiable text generation is challenging for the following reasons. Firstly, it often involves long text generation, where the focus of the content changes over time, characterized by focus-shifting phenomenon Lan and Jiang (2021); Sun et al. (2023a). This dynamic poses challenges in consistently aligning claims with the appropriate evidential references. For instance, as depicted in Figure 1, the discussion evolves from the health benefits of moderate coffee consumption to the adverse effects of excessive consumption. Such shifts demand a dynamic adaptation in the document pool to support the shifting focus of the generation. Secondly, identifying the intricate relationship between a claim and its potential evidence requires more than just linguistic matching—it demands careful analysis. For example, in Figure 1, the third claim suggests excessive coffee consumption leads to various mental and physical discomforts. The corresponding supporting document, although not explicitly mentioning these discomforts, describes symptoms that are inherently related to them. This necessitates an in-depth examination to confirm that the evidence truly supports the specific claim. Thirdly, striking a balance between the precision and breadth of retrieved documents presents a complex challenge for verifiable text generation. On the one hand, the task is susceptible to noisy documents during the claim-citation alignment process, emphasizing the need to retain a few highly relevant documents selectively. On the other hand, the intrinsic nature of verifiable text generation calls for a comprehensive collection of documents to enhance credibility. Therefore, crafting strategies to balance precision and breath in document retrieval is crucial for advancing verifiable text generation.
@@
When composing text with citations, individuals are capable of adaptively gathering the most relevant information regarding the claim being written, typically involving active information seeking and frequent verification. Inspired by this process, we propose Vtg, short for Verifiable Text Generation, a novel framework that operates through iterative generation and verification, utilizing an evolving memory and a two-tier verifier. Specifically, to address the challenge of focus-shifting, Vtg employs an evolving long short-term memory system. This system effectively archives important documents in long-term memory and maintains recent ones in short-term memory, thereby providing support for the evolving focus of the generation. Moreover, to identify the complex relationship between a claim and its potential evidence, Vtg employs a generation verifier and a memory verifier, both using Natural Language Inference (NLI) model to assess the logical support of potential evidence for the claim. The generation verifier first checks if the cited documents logically support the claim. If there’s a misalignment, the memory verifier reevaluates the claim against the documents stored in memory. A positive outcome suggests that the misalignment is due to the citation generation process, not because the information in the claim is wrong, leading to the adoption of a refined set of documents from memory for citation. Conversely, a negative outcome indicates potential factual inaccuracies in the claim, triggering an evidence finder to gather external information, which facilitates the regeneration of a more accurate and verifiable claim. Lastly, to balance between precision and breath in document retrieval, Vtg incorporates active retrieval and diverse query generation. Retrieval is initiated only when the claim does not pass the memory verifier, indicating potential factual inaccuracies. This approach guarantees the necessity of retrieval, reducing noise from unnecessary retrieval, and thereby enhancing retrieval precision. By instructing LLMs to generate diverse queries, the breath of retrieved documents is broadened, enabling the documents to offer comprehensive support for the claim.
@@ 3 Experiment
Correct Citation Correct Citation Correct Citation
Metrics EM F1 Rec Prec F1 EM Rec Prec F1 EM Rec Prec F1
@@
Vanilla 23.40 21.98 29.55 22.25 25.39 55.80 67.66 60.66 63.97 54.80 71.39 61.71 66.20
Summ 23.20 20.00 30.89 28.43 29.61 58.00 70.51 62.07 66.02 57.00 51.55 52.21 51.88
Snippet 21.80 20.05 25.18 21.95 23.45 58.40 53.44 49.15 51.21 57.20 43.56 41.43 42.47
Rerank 22.60 21.13 47.03 47.53 47.28 56.40 89.93 76.33 82.57 56.20 83.56 73.57 78.25
Vtg 25.60 23.27 55.36 49.59 52.32 60.00 92.16 86.51 89.25 58.00 88.69 82.02 85.22
@@
Vanilla 33.00 33.01 40.46 28.30 33.30 67.50 63.78 58.97 61.28 62.50 60.48 55.56 57.92
Summ 30.00 30.63 9.39 12.19 10.61 67.50 60.06 47.62 53.12 62.50 44.23 38.45 41.14
Snippet 32.00 30.13 13.86 18.49 15.84 67.00 65.41 52.32 58.14 62.00 54.72 46.99 50.56
Rerank 32.67 33.09 56.13 45.22 50.09 67.00 73.72 64.90 69.03 61.50 71.30 63.44 67.14
Vtg 41.50 40.19 63.89 57.65 60.61 68.00 93.00 88.72 90.81 63.00 91.85 86.59 89.14
@@
Correct Citation Correct Citation Correct Citation
Metrics EM D-F1 R-L DR Rec Prec F1 Claim Rec Prec F1 EM F1
@@
Vanilla 32.00 27.52 33.53 30.53 72.78 62.09 67.01 12.20 59.79 48.26 53.41 35.64 55.19
Summ 41.71 28.95 37.18 33.07 62.15 59.60 60.85 14.20 60.13 52.42 56.01 38.82 52.87
Snippet 39.22 27.01 35.65 31.33 46.23 47.04 46.63 14.33 31.47 32.72 32.08 38.19 39.17
Rerank 37.14 28.21 32.18 30.20 88.29 75.74 81.53 11.67 73.80 61.12 66.86 36.80 71.30
Vtg 41.92 30.53 37.87 34.20 89.15 82.57 85.73 14.73 81.50 72.16 76.55 40.05 77.81
@@
Vanilla 40.25 31.47 35.81 33.64 58.13 55.17 56.61 13.43 58.66 47.40 52.43 43.34 52.31
Summ 41.33 28.91 37.21 33.06 48.31 40.68 44.17 11.50 39.43 31.81 35.21 42.57 36.85
Snippet 39.60 30.11 38.35 34.23 53.14 43.19 47.65 13.67 45.29 37.23 40.87 42.85 42.61
Rerank 39.55 29.94 39.38 34.66 75.83 69.81 72.70 14.76 76.21 61.67 68.17 43.10 65.43
Vtg 41.53 31.64 39.45 35.55 86.70 79.95 83.19 16.67 82.63 71.56 76.70 46.14 80.09
@@ 3.1 Baselines
For an equitable comparison, we have selected the following four best-performing baseline methodologies as proposed in ALCE Gao et al. (2023b). For more details, please refer to Appendix A.
@@ 3.2 Datasets and Evaluation
In line with Jiang et al. (2023), LLMs are prompted to provide the final answer, which is then evaluated against the reference answer using answer-level exact match (EM), and token-level F1, precision, and recall metrics.
@@
For ASQA, we apply the metrics outlined in Stelmakh et al. (2022), including EM, a soft match using a RoBERTa-based QA model (Disambig-F1), ROUGE Lin (2004), and a combined DR score. In the case of ELI5, we adhere to the evaluation criteria of Gao et al. (2023b), focusing on whether the model’s predictions address the sub-claims of the gold-standard answer.
@@
Following the methodology of Yu et al. (2022) and Sun et al. (2023b), LLMs are prompted to generate the final answer, which is then compared with the reference answer using answer-level exact match (EM), and token-level F1.
Verifiability Evaluation
To evaluate the citation quality of responses, we employ the approach of Gao et al. (2023b), focusing on calculating Citation Recall, Citation Precision, and the combined Citation F1 score. Citation Recall examines whether the output is fully supported by the cited documents, while Citation Precision assesses the redundancy of the citations included.
@@ 3.3 Implementation Details
To prove the generalizability of our method, we conducted experiments using LLMs of different parameter sizes. Specifically, we utilized two LLMs: Vicuna-13B-v1.5-16k [^3] https://huggingface.co/lmsys/vicuna-13b-v1.5-16k Zheng et al. (2023) and Text-Davinci-003444https://api.openai.com/v1/completions as of October 2023 Ouyang et al. (2022) for evaluation, respectively. For the evaluation of verifiability and the inference tasks in both the Rerank and Vtg methods, we employed the TRUE [^5] https://huggingface.co/google/t5_xxl_true_nli_mixture model (Raffel et al., 2020), a T5-11B model fine-tuned on a collection of NLI datasets, to automatically examine whether the cited documents entail the claim. Following Gao et al. (2023b), we use Wikipedia dump from Dec. 20, 2018 as our retrieval corpus and use DPR Karpukhin et al. (2020) as our dense retriever.
@@ 3.4 Main Results
Third, the comparison between the two different LLMs reveals interesting findings. By integrating a broader range of documents, Summ and Snippet outperform Vanilla in terms of overall correctness when evaluated with Vicuna-13B. However, this advantage diminishes when evaluated with Text-Davinci-003. This could be attributed to Text-Davinci-003’s extensive internal knowledge base, which enables it to generate answers without relying on external sources, as evidenced by the performance comparison when applying the same method to both LLMs. Consequently, Summ and Snippet may introduce unnecessary noisy information in the context of Text-Davinci-003, leading to correctness degradation. Additionally, these methods struggle with citation quality on both LLMs, as simplified documents make it hard for LLMs to generate the correct citation.
@@ 3.5 Analyses
We examine the impact of retrieval parameters on Vtg’s performance using ASQA dataset, focusing on the number of generated queries $M$ and the retrieved document number per query $N$. As shown in Figure 3, we observe a consistent trend for both parameters. Initially, increasing $M$ and $N$ enhances both correctness and citation quality, which can be attributed to the fact that a larger pool of documents offers a broader knowledge scope and provides more citation options for the LLM. However, continually increasing them beyond a certain threshold results in a decline in performance, which is mainly because an excessively large document pool can introduce noisy information, negatively impacting both claim generation and citation generation processes.
@@
We analyze token consumption across various methods on the NQ dataset using Text-Davinci-003 as the LLM. As shown in Figure 4, Vanilla achieves the lowest token cost due to its single API request. After introducing more documents, Summ, Snippet incur higher token costs but with lower citation quality as simplified documents challenge LLMs’ citation generation. Rerank produces better citation quality than other baselines but incurs the highest token cost due to its strategy of multiple sampling. In contrast, our method Vtg demonstrates a lower token cost than Rerank while significantly improving citation quality, highlighting the superiority of our approach. Importantly, users can adjust Vtg’s token cost with the max trial parameter to balance performance and computational cost.

### v3

@@ Abstract
Towards Verifiable Text Generation with Evolving Memory
and Self-Reflection
Hao Sun1,2, Hengyi Cai3, Bo Wang5, Yingyan Hou3 Xiaochi Wei4, Shuaiqiang Wang4, Yan Zhang1,2, Dawei Yin4 1State Key Laboratory of General Artificial Intelligence, Peking University, Beijing, China 2School of Intelligence Science and Technology, Peking University 3Chinese Academy of Sciences, 4Baidu Inc, 5Beijing Institute of Technology sunhao@stu.pku.edu.cn
@@ 1 Introduction
However, verifiable text generation is challenging for the following reasons. Firstly, it often involves long text generation, where the focus of the content changes over time, characterized by focus-shifting phenomenon Lan and Jiang (2021); Sun et al. (2023a). This dynamic poses challenges in consistently aligning claims with the appropriate evidential references. For instance, as depicted in Figure 1, the discussion evolves from the health benefits of moderate coffee consumption to the adverse effects of excessive consumption. Such shifts demand a dynamic adaptation in the document pool to support the shifting focus of the generation. Secondly, identifying the intricate relationship between a claim and its potential evidence requires more than just linguistic matching—it demands careful analysis. For example, in Figure 1, the third claim suggests excessive coffee consumption leads to various mental and physical discomforts. The corresponding supporting document, although not explicitly mentioning these discomforts, describes symptoms inherently related to them. This necessitates an in-depth examination to confirm that the evidence truly supports the specific claim. Thirdly, striking a balance between the precision and breadth of retrieved documents presents a complex challenge for verifiable text generation. On the one hand, the task is susceptible to noisy documents during the claim-citation alignment process, emphasizing the need to selectively retain a few highly relevant documents. On the other hand, the intrinsic nature of verifiable text generation calls for a comprehensive collection of documents to enhance credibility. Therefore, crafting strategies to balance precision and breadth in document retrieval is crucial for advancing verifiable text generation.
@@
When composing text with citations, individuals are capable of adaptively gathering the most relevant information regarding the claim being written, typically involving active information seeking and frequent verification. Inspired by this process, we propose Vtg, short for Verifiable Text Generation, a novel framework that operates through iterative generation and verification, utilizing an evolving memory and a two-tier verifier. Specifically, to address the challenge of focus-shifting, Vtg employs an evolving long short-term memory system. This system effectively archives important documents in long-term memory and maintains recent ones in short-term memory, thereby providing support for the evolving focus of the generation. Moreover, to identify the complex relationship between a claim and its potential evidence, Vtg employs a generation verifier and a memory verifier, both using Natural Language Inference (NLI) model to assess the logical support of potential evidence for the claim. The generation verifier first checks if the cited documents logically support the claim. If there’s a misalignment, the memory verifier reevaluates the claim against the documents stored in memory. A positive outcome suggests that the misalignment is due to the citation generation process, not because the information in the claim is wrong, leading to the adoption of a refined set of documents from memory for citation. Conversely, a negative outcome indicates potential factual inaccuracies in the claim, triggering an evidence finder to gather external information, which facilitates the regeneration of a more accurate and verifiable claim. Lastly, to balance between precision and breadth in document retrieval, Vtg incorporates active retrieval and diverse query generation. Retrieval is initiated only when the claim does not pass the memory verifier, indicating potential factual inaccuracies. This approach guarantees the necessity of retrieval, reducing noise from unnecessary retrieval, and thereby enhancing retrieval precision. By instructing LLMs to generate diverse queries, the breadth of retrieved documents is broadened, enabling the documents to offer comprehensive support for the claim.
@@ 3 Experiment
Correct ALCE.Citation LLM.Citation Correct ALCE.Citation LLM.Citation Correct ALCE.Citation LLM.Citation
Metrics EM F1 Rec Prec F1 Rec Prec F1 EM Rec Prec F1 Rec Prec F1 EM Rec Prec F1 Rec Prec F1
@@
Vanilla 23.40 21.98 29.55 22.25 25.39 41.59 35.04 38.03 55.80 67.66 60.66 63.97 67.50 67.83 67.67 54.80 71.39 61.71 66.20 77.46 63.94 70.05
Summ 23.20 20.00 30.89 28.43 29.61 37.66 39.12 38.37 58.00 70.51 62.07 66.02 68.23 66.57 67.39 57.00 51.55 52.21 51.88 56.76 62.86 59.66
Snippet 21.80 20.05 25.18 21.95 23.45 33.67 29.75 31.59 58.40 53.44 49.15 51.21 68.46 69.57 69.01 57.20 43.56 41.43 42.47 57.56 59.57 58.55
Rerank 22.60 21.13 47.03 47.53 47.28 54.33 53.73 54.03 56.40 89.93 76.33 82.57 88.20 67.77 76.64 56.20 83.56 73.57 78.25 81.66 73.49 77.36
Vtg 25.60 23.27 55.36 49.59 52.32 62.76 54.69 58.45 60.00 92.16 86.51 89.25 89.43 81.38 85.21 58.00 88.69 82.02 85.22 86.35 78.06 82.00
@@
Vanilla 33.00 33.01 40.46 28.30 33.30 59.07 43.00 49.77 67.50 63.78 58.97 61.28 71.18 66.52 68.77 62.50 60.48 55.56 57.92 66.45 61.59 63.93
Summ 30.00 30.63 9.39 12.19 10.61 23.19 24.64 23.89 67.50 60.06 47.62 53.12 68.33 56.65 61.95 62.50 44.23 38.45 41.14 55.36 49.20 52.10
Snippet 32.00 30.13 13.86 18.49 15.84 37.36 38.99 38.16 67.00 65.41 52.32 58.14 71.81 68.07 69.89 62.00 54.72 46.99 50.56 73.05 69.55 71.25
Rerank 32.67 33.09 56.13 45.22 50.09 63.43 46.32 53.54 67.00 73.72 64.90 69.03 78.12 70.00 73.84 61.50 71.30 63.44 67.14 79.03 66.37 72.15
Vtg 41.50 40.19 63.89 57.65 60.61 70.47 59.13 64.30 68.00 93.00 88.72 90.81 90.70 87.52 89.08 63.00 91.85 86.59 89.14 84.92 73.20 78.63
@@
Correct ALCE.Citation LLM.Citation Correct ALCE.Citation LLM.Citation Correct Citation
Metrics EM D-F1 R-L DR Rec Prec F1 Rec Prec F1 Claim Rec Prec F1 Rec Prec F1 EM F1
@@
Vanilla 32.00 27.52 33.53 30.53 72.78 62.09 67.01 73.28 66.59 69.78 12.20 59.79 48.26 53.41 81.46 76.89 79.11 35.64 60.06
Summ 41.71 28.95 37.18 33.07 62.15 59.60 60.85 68.95 70.19 69.56 14.20 60.13 52.42 56.01 77.87 72.42 75.05 38.82 57.44
Snippet 39.22 27.01 35.65 31.33 46.23 47.04 46.63 56.55 63.03 59.61 14.33 31.47 32.72 32.08 46.69 49.32 47.97 38.19 46.26
Rerank 37.14 28.21 32.18 30.20 88.29 75.74 81.53 88.29 75.74 81.53 11.67 73.80 61.12 66.86 84.57 77.09 80.65 36.80 72.67
Vtg 41.92 30.53 37.87 34.20 89.15 82.57 85.73 89.15 82.57 85.73 14.73 81.50 72.16 76.55 87.60 84.46 86.00 40.05 78.65
@@
Vanilla 40.25 31.47 35.81 33.64 58.13 55.17 56.61 58.13 55.17 56.61 13.43 58.66 47.40 52.43 58.66 47.40 52.43 43.34 55.31
Summ 41.33 28.91 37.21 33.06 48.31 40.68 44.17 50.48 44.44 47.27 11.50 39.43 31.81 35.21 52.27 48.47 50.30 42.57 41.98
Snippet 39.60 30.11 38.35 34.23 53.14 43.19 47.65 59.31 52.05 55.44 13.67 45.29 37.23 40.87 62.39 55.19 58.57 42.85 50.64
Rerank 39.55 29.94 39.38 34.66 75.83 69.81 72.70 76.41 70.01 73.07 14.76 76.21 61.67 68.17 86.98 77.64 82.04 43.10 68.18
Vtg 41.53 31.64 39.45 35.55 86.70 79.95 83.19 89.10 79.84 84.22 16.67 82.63 71.56 76.70 87.94 81.79 84.75 46.14 80.14
@@ 3.1 Baselines
For an equitable comparison, we have selected the following four best-performing baseline methodologies as proposed in ALCE Gao et al. (2023b).
@@
Besides, we also compare with several post-processing baselines, which include PostCite, RefineCite, VeriCite and VeriRefine. Due to space limits, we only put the four baselines from ALCE in the main results, for the complete experiment results, please refer to Appendix A.
@@ 3.2 Datasets and Evaluation
In line with Jiang et al. (2023), LLMs are prompted to provide the final answer, which is then evaluated against the reference answer using answer-level Exact Match (EM), token-level precision, recall and F1 metrics.
@@
For ASQA, we apply the metrics outlined in Stelmakh et al. (2022), including Exact Match (EM), a soft match using a RoBERTa-based QA model (Disambig-F1), ROUGE Lin (2004) and a combined DR score. In the case of ELI5, we adhere to the evaluation criteria of Gao et al. (2023b), focusing on whether the model’s predictions address the sub-claims of the gold-standard answer.
@@
Following the methodology of Yu et al. (2022) and Sun et al. (2023b), LLMs are prompted to generate the final answer, which is then compared with the reference answer using answer-level Exact Match (EM).
Verifiability Evaluation.
To evaluate the citation quality of responses, we employ the approach of Gao et al. (2023b), focusing on calculating ALCE.Citation Recall, ALCE.Citation Precision, and the combined ALCE.Citation F1 score. ALCE.Citation Recall examines whether the output is fully supported by the cited documents, while ALCE.Citation Precision assesses the redundancy of the citations included. Additionally, to further improve the robustness of the evaluation, we incorporate the use of LLM as the citation evaluator. Specifically, The Qwen-Max is given a sentence and all the passages that the sentence cited and is asked to judge whether the passages fully support the sentence, which is used as the signal to compute LLM.Citation Recall and LLM.Citation Precision.
@@ 3.3 Implementation Details
To prove the generalizability of our method, we conduct experiments using LLMs of different parameter sizes. Specifically, we utilize two LLMs: Vicuna-13B-v1.5-16k [^3] https://huggingface.co/lmsys/vicuna-13b-v1.5-16k Zheng et al. (2023) and Text-Davinci-003444https://api.openai.com/v1/completions as of October 2023 Ouyang et al. (2022) for evaluation, respectively. For the evaluation of verifiability and the inference tasks in both the Rerank and Vtg methods, we employ the TRUE [^5] https://huggingface.co/google/t5_xxl_true_nli_mixture model (Raffel et al., 2020), a T5-11B model fine-tuned on a collection of NLI datasets, to automatically examine whether the cited documents entail the claim. We also experiment with different NLI models for inference and evaluation, please refer to Appendix D for more details. Following Gao et al. (2023b), we use Wikipedia dump from Dec. 20, 2018 as our retrieval corpus and use DPR Karpukhin et al. (2020) as our dense retriever.
@@ 3.4 Main Results
Third, the comparison between the two different LLMs reveals interesting findings. By integrating a broader range of documents, Summ and Snippet outperform Vanilla in terms of overall correctness when evaluated with Vicuna-13B. However, this advantage diminishes when evaluated with Text-Davinci-003. This could be attributed to Text-Davinci-003’s extensive internal knowledge base, which enables it to generate answers without relying on external sources, as evidenced by the performance comparison when applying the same method to both LLMs. Consequently, Summ and Snippet may introduce unnecessary noisy information in the context of Text-Davinci-003, leading to correctness degradation. Additionally, these methods struggle with citation quality on both LLMs, as simplified documents make it hard for LLMs to generate the correct citations.
@@ 3.5 Analysis
We examine the impact of retrieval parameters on Vtg’s performance using ASQA dataset, focusing on the number of generated queries $M$ and the number of retrieved documents per query $N$. As shown in Figure 3, we observe a consistent trend for both parameters. Initially, increasing $M$ and $N$ enhances both correctness and citation quality, which can be attributed to the fact that a larger pool of documents offers a broader knowledge scope and provides more citation options for the LLM. However, continually increasing them beyond a certain threshold results in a decline in performance, which is mainly because an excessively large document pool can introduce noisy information, negatively impacting both claim generation and citation generation processes.
@@
We analyze token consumption across various methods on the NQ dataset using Text-Davinci-003 as the LLM. As shown in Figure 4, Vanilla achieves the lowest token cost due to its single API request. After introducing more documents, Summ, Snippet incur higher token costs but with lower citation quality as simplified documents challenge LLMs’ citation generation. Rerank produces better citation quality than other baselines but incurs the highest token cost due to its strategy of multiple sampling. In contrast, our method Vtg demonstrates a lower token cost than Rerank while significantly improving citation quality, highlighting the superiority of our approach. Importantly, users can adjust Vtg’s token cost with the max trial parameter $T$ to balance performance and computational cost.
@@ 5 Conclusion
Limitations
In this work, we propose a novel frame Vtg for verifiable text generation. The limitations of the proposed method are as follows: (1) the computational cost of Vtg is relatively high due to the need for multiple API calls and frequent verification. This may restrict its applicability in resource-intensive scenarios or systems with limited computational resources; (2) the effectiveness of our verification process is constrained by the precision of the NLI models. In instances where the NLI model’s accuracy is suboptimal, there is a risk of incorporating erroneous information into the method, potentially compromising the verifiability of the generated text.
As for future work, we plan to mitigate the computational cost of the method by developing more efficient pipelines. Moreover, we aim to reduce our approach’s reliance on NLI models, thereby enhancing the overall robustness of our framework.
Acknowledgement
This work is supported in part by Ucap Cloud and the State Key Laboratory of General Artificial Intelligence.
Ethics Statement
This work was conducted in rigorous compliance with the ACL Ethics Policy. All datasets and large language models (LLMs) used for evaluation are publicly available. Furthermore, our work strives to improve the verifiability of LLM outputs, which could potentially broaden the application scenarios of LLMs. We do not foresee any form of negative ethical impact induced by our work.
@@ Appendix A Baselines
PostCite:
The LLM first generates answers then cites the most relevant passage from the top-100 retrieved documents for each statement.
RefineCite:
The LLM first produces answers with citations, then uses NLI to remove unnecessary citations, refining the citation set.
VeriCite:
The LLM first generates sentences with citations and then uses an NLI method to ensure the citation set conclusively supports the sentence. If not, it finds and cites the best matching passage from the top-100 retrieved documents using GTR.
VeriRefine:
Integrates RefineCite’s refinement and VeriCite’s verification processes, ensuring citations are both necessary and fully supportive, optimizing citation accuracy and relevance
Datasets Wikihop WebQ NQ
Correct ALCE.Citation LLM.Citation Correct ALCE.Citation LLM.Citation Correct ALCE.Citation LLM.Citation
Metrics EM F1 Rec Prec F1 Rec Prec F1 EM Rec Prec F1 Rec Prec F1 EM Rec Prec F1 Rec Prec F1
Vicuna-13B
Vanilla 23.40 21.98 29.55 22.25 25.39 41.59 35.04 38.03 55.80 67.66 60.66 63.97 67.50 67.83 67.67 54.80 71.39 61.71 66.20 77.46 63.94 70.05
Summ 23.20 20.00 30.89 28.43 29.61 37.66 39.12 38.37 58.00 70.51 62.07 66.02 68.23 66.57 67.39 57.00 51.55 52.21 51.88 56.76 62.86 59.66
Snippet 21.80 20.05 25.18 21.95 23.45 33.67 29.75 31.59 58.40 53.44 49.15 51.21 68.46 69.57 69.01 57.20 43.56 41.43 42.47 57.56 59.57 58.55
Rerank 22.60 21.13 47.03 47.53 47.28 54.33 53.73 54.03 56.40 89.93 76.33 82.57 88.20 67.77 76.64 56.20 83.56 73.57 78.25 81.66 73.49 77.36
PostCite 22.20 15.01 8.34 6.90 7.55 22.59 15.76 18.57 58.60 46.26 32.73 38.34 65.66 61.11 63.30 46.60 35.21 24.40 28.83 52.88 50.98 51.91
RefineCite 24.20 22.15 48.00 39.50 43.34 52.61 47.12 49.72 55.40 78.04 72.55 75.19 84.44 79.72 82.01 55.20 65.05 61.80 63.39 75.47 77.00 76.23
VeriCite 22.89 21.29 14.69 8.21 10.53 41.36 24.77 30.98 57.20 80.97 54.41 65.08 78.97 69.07 73.69 54.80 75.81 54.17 63.19 83.81 65.83 73.74
VeriRefine 24.61 23.06 15.90 11.30 13.21 47.57 32.97 38.94 54.40 82.33 64.67 72.44 81.33 75.00 78.04 55.60 71.34 58.35 64.19 81.34 67.52 73.79
Vtg 25.60 23.27 55.36 49.59 52.32 62.76 54.69 58.45 60.00 92.16 86.51 89.25 89.43 81.38 85.21 58.00 88.69 82.02 85.22 86.35 78.06 82.00
Text-Davinci-003
Vanilla 33.00 33.01 40.46 28.30 33.30 59.07 43.00 49.77 67.50 63.78 58.97 61.28 71.18 66.52 68.77 62.50 60.48 55.56 57.92 66.45 61.59 63.93
Summ 30.00 30.63 9.39 12.19 10.61 23.19 24.64 23.89 67.50 60.06 47.62 53.12 68.33 56.65 61.95 62.50 44.23 38.45 41.14 55.36 49.20 52.10
Snippet 32.00 30.13 13.86 18.49 15.84 37.36 38.99 38.16 67.00 65.41 52.32 58.14 71.81 68.07 69.89 62.00 54.72 46.99 50.56 73.05 69.55 71.25
Rerank 32.67 33.09 56.13 45.22 50.09 63.43 46.32 53.54 67.00 73.72 64.90 69.03 78.12 70.00 73.84 61.50 71.30 63.44 67.14 79.03 66.37 72.15
Vtg 41.50 40.19 63.89 57.65 60.61 70.47 59.13 64.30 68.00 93.00 88.72 90.81 90.70 87.52 89.08 63.00 91.85 86.59 89.14 84.92 73.20 78.63
Datasets ASQA ELI5 Overall
Correct ALCE.Citation LLM.Citation Correct ALCE.Citation LLM.Citation Correct Citation
Metrics EM D-F1 R-L DR Rec Prec F1 Rec Prec F1 Claim Rec Prec F1 Rec Prec F1 EM F1
Vicuna-13B
Vanilla 32.00 27.52 33.53 30.53 72.78 62.09 67.01 73.28 66.59 69.78 12.20 59.79 48.26 53.41 81.46 76.89 79.11 35.64 60.06
Summ 41.71 28.95 37.18 33.07 62.15 59.60 60.85 68.95 70.19 69.56 14.20 60.13 52.42 56.01 77.87 72.42 75.05 38.82 57.44
Snippet 39.22 27.01 35.65 31.33 46.23 47.04 46.63 56.55 63.03 59.61 14.33 31.47 32.72 32.08 46.69 49.32 47.97 38.19 46.26
Rerank 37.14 28.21 32.18 30.20 88.29 75.74 81.53 88.29 75.74 81.53 11.67 73.80 61.12 66.86 84.57 77.09 80.65 36.80 72.67
PostCite 25.55 22.24 35.36 28.80 38.54 25.69 30.83 47.14 38.14 42.16 14.47 26.90 17.02 20.85 62.15 52.20 56.74 33.48 35.91
RefineCite 36.30 28.24 35.35 31.79 77.80 60.50 68.07 69.80 61.17 65.20 11.60 26.90 59.05 36.96 32.15 67.46 43.54 36.54 60.36
VeriCite 34.36 28.86 36.16 32.51 76.15 72.67 74.37 77.32 75.83 76.57 13.73 64.93 40.14 49.61 77.93 62.48 69.35 36.60 58.71
VeriRefine 34.36 28.86 35.40 32.13 80.80 59.28 68.39 76.80 64.95 70.38 11.27 61.57 37.59 46.68 64.57 51.26 57.15 36.05 58.32
Vtg 41.92 30.53 37.87 34.20 89.15 82.57 85.73 89.15 82.57 85.73 14.73 81.50 72.16 76.55 87.60 84.46 86.00 40.05 78.65
Text-Davinci-003
Vanilla 40.25 31.47 35.81 33.64 58.13 55.17 56.61 58.13 55.17 56.61 13.43 58.66 47.40 52.43 58.66 47.40 52.43 43.34 55.31
Summ 41.33 28.91 37.21 33.06 48.31 40.68 44.17 50.48 44.44 47.27 11.50 39.43 31.81 35.21 52.27 48.47 50.30 42.57 41.98
Snippet 39.60 30.11 38.35 34.23 53.14 43.19 47.65 59.31 52.05 55.44 13.67 45.29 37.23 40.87 62.39 55.19 58.57 42.85 50.64
Rerank 39.55 29.94 39.38 34.66 75.83 69.81 72.70 76.41 70.01 73.07 14.76 76.21 61.67 68.17 86.98 77.64 82.04 43.10 68.18
Vtg 41.53 31.64 39.45 35.55 86.70 79.95 83.19 89.10 79.84 84.22 16.67 82.63 71.56 76.70 87.94 81.79 84.75 46.14 80.14
The complete experimental result is shown in Table 8 and Table 9.